OpenStudy (anonymous):

Find 2 numbers whose sum is 64 and whose difference is 42.

OpenStudy (anonymous):

i got 53,11

OpenStudy (anonymous):

Yes you are absolutely right..

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

n1 + n2 =64 |n1 - n2| = 42 from 1º equation: n1 = 64 -n2 put it in 2º equation: |64-n2 -n2| = 42 solve for n2: |64 -2n2| = 42 so 2 posibilities: 64-2n2=-42 and 64-2n2=42 106=2n2 n2 = 53 and 64 -42 =2n2 2n2=22 n2=11

OpenStudy (anonymous):

k